OpenStudy (anonymous):

How do you find the period of a sine or cosine function? Example: .5sin(1/6)x

OpenStudy (jdoe0001):

\(\bf y = {\color{purple}{ A}} ( {\color{blue}{ B}}x + {\color{red}{ C}} ) + {\color{green}{ D}} \\ \quad \\ \quad \\ period=\cfrac{2\pi }{{\color{blue}{ B}}}\)

OpenStudy (jdoe0001):

well... in this case I assume \(\bf y = {\color{purple}{ 5}} sin \left( {\color{blue}{ \frac{1}{6}}}x + {\color{red}{ 0}} \right) + {\color{green}{ 0}} \\ \quad \\ \quad \\ period=\cfrac{2\pi }{{\color{blue}{ B}}}\)